Circular No. 5609

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

V645 CENTAURI G. F. Benedict, McDonald Observatory, reports for the Hubble Space

Telescope Astrometry Science Team: “We announce the discovery of a nearly-periodic variation in the brightness of V645 (Proxima) Cen. Science Verification and GTO astrometry observations secured with HST Fine Guidance Sensor #3 over the past 154 days indicate a period of 42 ± 1 days with a half-amplitude of 0.012 ± 0.001 magnitude. The FGS bandpass used to obtain these data has FWHM = 234 nm, centered at 583 nm. Data acquisition will continue with HST and a ten-day sampling rate through 1993 Apr., with a gap from roughly 1992 Nov. 1 to Dec. 20, due to HST sunpointing constraints. There is evidence that we have undersampled the light curve. We encourage broadband photometric monitoring of Proxima Cen with higher temporal resolution than presently scheduled with HST.”

DRISSEN V211b L. Drissen and M. Shara, Space Telescope Science Institute; and M.

Dopita and D. Wickramasinghe, Australian National University (ANU), report: “We have discovered a new AM Her-type variable, based on a survey of the overlap regions of U.K. Schmidt plates for large-amplitude variables. The object, Drissen V211b, of apparent magnitude V ∼ 18, is located at α = 20^{h}04^{m}17^{s}.86, δ = −65^{o}36′08′′.0 (equinox 1950.0). A spectrogram obtained at the Anglo-Australian Telescope shows the strong emission lines of He II and H I typical of the class. A period of 2^{h}40^{m} was measured by M. Dopita and J. Bell (ANU) using CCD photometry on the ANU 2.3-m telescope at Siding Spring. Both the period and the AM Her nature of the object was confirmed by photometric, polarimetric, and spectroscopic measurements by J. Bailey (Anglo-Australian Observatory), D. Wickramasinghe (ANU), and J. Hough (University of Hertfordshire) using the Hatfield polarimeter and the AAO spectropolarimeter on the AAT. The continuum polarization ranges from –20 to +5 percent. The full amplitude of the velocity variation about the orbit is 650 km/s. This variable is of unique interest, being of AM Her type in the period gap of cataclysmic variables.”

FG SAGITTAE Representative visual magnitude estimates (cf. IAUC 5604), from E.

Waagen, AAVSO: Sept. 5.16 UT, 10.5 (R. Royer, Lakewood, CA); 9.07, 10.8 (P. Sventek, Houston, TX); 10.17, 10.7 (Royer).

1992 September 11 Daniel W. E. Green